Let ?? be the probability that in a group of ? people, at least two share the same birthday.

Assume there are 365 days in a year, and that all birthdays are equally likely.

a)

What is the probability that in a group of 2, 3, 4, or 5 people, at least two have the same birthday?

?2=?

?3=?

?4=?

?5=?
Answer #1

P(2) =P(at least two have the same birthday) =1-P(both have differnt B'day)=(365*364)/3652 =0.002740

P(3)=P(at least two have the same birthday) =1-P(both have differnt B'day)=(365*364*363)/3653 =0.008204

P(4)==(365*364*363*362)/3654 =0.016356

P(5)==0.027136